OpenStudy (yuliacosta111):

Joan is going to pick up Gary and Kash for band practice at her house this afternoon. She is going to pick up Kash first, then Gary, and then go back home for practice. How many total miles will Joan travel while picking up her bandmates? Round your answer to the nearest mile. 9 miles 10 miles 19 miles 22 miles

OpenStudy (yuliacosta111):

the graph shows the location of Gary's, Joan's, and Kash's houses on a coordinate plane. Each unit on the graph represents 1 mile.

OpenStudy (yuliacosta111):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

first of all lets find way Joan-Kash

OpenStudy (alexandervonhumboldt2):

distance formula: \[\huge\mathbb {d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}}\]

OpenStudy (alexandervonhumboldt2):

what do we get?

OpenStudy (yuliacosta111):

i think 8

OpenStudy (alexandervonhumboldt2):

8 miles total? that is impossible.

OpenStudy (alexandervonhumboldt2):

the one only distance from joan to kash is 9.43 miles. Now count by grid the other distances: we get that kash-gary=8 and gary-joan=5 thus 8+5+9.43=?

OpenStudy (yuliacosta111):

22.43

OpenStudy (alexandervonhumboldt2):

yes. now in rounded to nearest mile 22.43 is ?

OpenStudy (yuliacosta111):

22

OpenStudy (alexandervonhumboldt2):

yeah